OpenStudy (anonymous):

You buy new shoes at Walmart early one morning, The shoes you want are 30% off the ticket price and for today only there is an additional 15 percent off for early birds. After calculating the 7.5 percent tax, how much do you end up spending?

OpenStudy (anonymous):

@rajathsbhat

OpenStudy (unklerhaukus):

how much do the shoes cost ?

OpenStudy (anonymous):

Please note it does not mention dollar amounts. :/

OpenStudy (anonymous):

This is why i was confused.

OpenStudy (anonymous):

x*0.55*1,075=answer since you get 45% discount + 7,5 tax

OpenStudy (anonymous):

so, do i do x*0.55*,1,075?

OpenStudy (anonymous):

where x is the price.

OpenStudy (anonymous):

ok, I solve for x.

OpenStudy (anonymous):

Will do, one moment please.

OpenStudy (anonymous):

Hmmm. I cannot solve it with my calculator.

OpenStudy (anonymous):

Just 0.55*1,075 maybe?

OpenStudy (anonymous):

raja???

OpenStudy (anonymous):

I did that and got \[x= 473/800\]

OpenStudy (anonymous):

No, you got (473/800)*x=what you spend. NOT equal x

OpenStudy (anonymous):

Oh.

OpenStudy (anonymous):

Umm. Have you gotten an answer?

OpenStudy (anonymous):

From what you asked, thats the answer (473/800)*x=y Where x is the initial price and y= what you spend.

OpenStudy (anonymous):

Oh, so I write that as the answer, as it is?

OpenStudy (anonymous):

just 473/800*x=y?

OpenStudy (anonymous):

yepp, but then you should explain what x and y stands for.

OpenStudy (anonymous):

Okedoke. Thanks!